OpenStudy (anonymous):

Help Please! Been suck for an hour now... Find the function with the given derivative whose graph passes though the point P. g'(x)=(1/x^2)+2x , P(-1,1)

OpenStudy (amistre64):

integrate the function ... you might know it as antiderivative?

OpenStudy (amistre64):

\[\int f(x)~dx\to F(x)+C\]

OpenStudy (anonymous):

How do you anitderive the function?

OpenStudy (amistre64):

as far as functions go; this is a pretty tame function. What do you know about an antiderivative? or derivatives for that matter?

OpenStudy (anonymous):

i know that an anitderivative is a function of f(x) whose derivative passes though a given point?

OpenStudy (amistre64):

you should have gone over some rules of derivatives ... i see one rule that would be most beneficial here. The power rule.\[\frac{d}{dx}[x^n]=n~x^{n-1}\]

OpenStudy (anonymous):

Oh.. I aware of the rules, but I have a hard time applying them...

OpenStudy (anonymous):

I'm*

OpenStudy (amistre64):

the antiderivative of this is working it backwards .... \[\int n~x^{n-1}~dx\to\frac{n}{n-1+1}x^{n-1+1}+C\]\[\frac{n}{n}x^{n}+C\]\[x^{n}+C\]

OpenStudy (amistre64):

add one to the exponent and divide it out ...

OpenStudy (anonymous):

I haven't learned integrals yet.. is there a way you could show it without the

OpenStudy (amistre64):

\[\frac{1}{x^2}+2x \] \[x^{-2}+2x \] \[\frac{1}{-2+1}x^{-2+1}+\frac{1}{1+1}2x^{1+1}+C \]

OpenStudy (amistre64):

there is ot symbol for "antiderivative", so not really ...

OpenStudy (anonymous):

okay. but why are powers adding? dont they minus?

OpenStudy (amistre64):

you are UNDOING a derivative ... so no, they dont minus. In order to undo multiplication, we divide. In order to undo subtraction, we add.

OpenStudy (anonymous):

ohhh i see. okay.

OpenStudy (amistre64):

\[\frac d{dx}[x^2]\to2x^{2-1}\] \[\int 2x^{2-1}~dx\to\frac{2}{2-1+1}x^{2-1+1}+C=x^2+C\]

OpenStudy (anonymous):

wow, that was a huge step.. haha okay.

OpenStudy (amistre64):

when we do this "undoing" power rule (an inverse operation) on your content, we get: \[\int~x^{-2}+2x~dx\to~\frac{1}{-2+1}x^{-2+1}+\frac{1}{1+1}2x^{1+1}+C\] \[-x^{-1}+x^2+C\]

OpenStudy (amistre64):

you are given a point in order to find a suitable value for C P(-1,1) \[1 = (-1)^{-1}+(-1)^2+C\] \[1 = -1+1+C\] \[1 = C\]

OpenStudy (anonymous):

so the function is \[-x^-1+x^2+C\]

OpenStudy (anonymous):

and we evaluated it at the point (-1,1) X=-1 which was just plugged into the function?

OpenStudy (anonymous):

@amister64

OpenStudy (amistre64):

yes \[g'(x) \to G(x)+C\]then evaluate it at the given point to determine the appropriate value of C

OpenStudy (amistre64):

and of course you have to read thru my typos :) i dropped a negative on accident when I solved for C

OpenStudy (anonymous):

I believe it was correct though right? C=1? what negative are you talking about that you dropped?

OpenStudy (amistre64):

\[that~one \to\color{red}{(-)}x^{-1}+x^2+C\] \[1=-(-1)^{-1}+(-1)^2+C\] \[1=1+1+C\] \[-1=C\]

OpenStudy (anonymous):

Ohhhhh!! okay. I was looking for your mistake but i couldn't find it haha. so C=-1!

OpenStudy (amistre64):

:)

OpenStudy (anonymous):

Thank you so much! I really appreciate all the help! (:

OpenStudy (amistre64):

youre welcome, and good luck ;)